Evaluate the functions when x = -2, 0, and 3
1)f(x) = -0.75x - 1
2)d(x) = -3/2x + 5

Answer by Alan3354(69443)

Evaluate the functions when x = -2, 0, and 3
1)f(x) = -0.75x - 1
--------
To find f(-2) sub -2 wherever there's an x
f(-2) = -0.75*(-2) - 1
f(-2) = 1.5 - 1
f(-2) = 0.5
=============
Do the others the same way.
--------------------------------------
2)d(x) = -3/2x + 5